# Convex cocompact groups in real hyperbolic spaces with limit set a Pontryagin sphere

### v2

@@ 1. Introduction
The Sierpiński carpet is the fractal curve obtained from a square by removing smaller and smaller squares, similarly to the way one obtains the Cantor set from an interval. In particular, the Sierpiński carpet has a well-defined collection of boundary squares. The Pontryagin sphere is obtained from the Sierpiński carpet by identifying the opposite sides of each boundary square as one would glue the edges of a square to form a torus. This results in a $2$-dimensional compact metrizable space without non-empty open subsets embeddable in $\mathbb{R}^{2}$; see Figure 1. The Pontryagin sphere can also be described as the limit of a certain inverse system of closed connected orientable surfaces of positive genus; see [JA91, FI03, ZA10, SW20, SW20sp] and §2.2. While this space is known to appear as the Gromov boundary of many Gromov hyperbolic groups, here we consider the problem of realizing the Pontryagin sphere as the limit set of concrete reflection groups in the isometry group $\mathrm{Isom}(\mathbb{H}^{d})$ of a real hyperbolic space $\mathbb{H}^{d}$.
@@
The starting point of our work is the following. If $W$ is a Gromov hyperbolic right-angled Coxeter group (RACG) whose nerve is a closed connected orientable surface of positive genus then the Gromov boundary $\partial_{\infty}W$ of $W$ is a Pontryagin sphere; see [FI03]. Hence, the up-shot of this paper is to find convex cocompact actions of two such Coxeter groups on $\mathbb{H}^{4}$ and $\mathbb{H}^{6}$, respectively.
@@
The Menger curve is a fractal curve that is obtained from a cube by removing smaller and smaller cubes, similarly to the way one obtains the Sierpiński carpet from a square and the Cantor set from an interval. It can be characterized as the unique compact, metrizable, connected, locally connected, $1$-dimensional space without local cut points and without non-empty open subsets embeddable in $\mathbb{R}^{2}$; see [AN58, AN58b]. According to [KP96], the Pontryagin sphere has a dense collection of Menger curves, so it is natural to look for Menger curves that are limit sets of quasiconvex subgroups of $\Gamma^{4}$ or $\Gamma^{6}$. In Theorem 5.3 we show that both $\Gamma^{4}$ and $\Gamma^{6}$ have convex cocompact subgroups with limit set homeomorphic to the Menger curve. These subgroups are obtained by forgetting one (or more, suitably chosen) reflections. Previous examples of convex cocompact subgroups of $\mathrm{Isom}(\mathbb{H}^{4})$ with Menger curve limit set were constructed by Bourdon [BO97], see Remark 5.4.
@@
Note that the analogous statement for Gromov hyperbolic groups with boundary a $2$-sphere, namely, that every group of the latter form possesses a quasiconvex subgroup with boundary a Sierpiński carpet, would follow from the Cannon conjecture, but even this implication requires non-trivial results about hyperbolic $3$-manifolds, including several of the ingredients that went into the solution to the virtual Haken conjecture [Ber15].
@@
We conclude with a few comments on realizing the Pontryagin sphere as a limit set in other spaces. By embedding $\mathbb{H}^{4}$ as a totally geodesic subspace in higher-dimensional hyperbolic spaces, one can realize $\Gamma^{4}$ as a convex cocompact subgroup of $\mathrm{Isom}(\mathbb{H}^{d})$ for each $d\geqslant 4$. It is then not difficult to show that, for each $d\geqslant 4$, there is a finite-index reflection subgroup of $\Gamma^{4}$ that admits arbitrarily small deformations that are Zariski dense in $\mathrm{Isom}(\mathbb{H}^{d})$ (note that sufficiently small deformations will still be convex cocompact reflection groups). As for the complex hyperbolic spaces $\mathbb{H}^{d}_{\mathbb{C}}$, it does not appear to be known whether there is a convex cocompact subgroup of $\mathrm{Isom}(\mathbb{H}_{\mathbb{C}}^{2})$ with limit set a Pontryagin sphere, though Granier [Granier] constructed convex cocompact subgroups of $\mathrm{Isom}(\mathbb{H}_{\mathbb{C}}^{2})$ with limit set a Menger curve (see also recent related work of Ma–Xie [MaXie]). Jeffrey Danciger and Theodore Weisman have communicated to us that they can show that the deformations of certain cusped hyperbolic $3$-orbifold groups inside $\mathrm{Isom}(\mathbb{H}_{\mathbb{C}}^{3})$ constructed in [PT23] descend to convex cocompact representations of suitable quotients with limit set a Pontryagin sphere. Outside of the world of symmetric spaces, one can use Ontaneda’s Riemannian hyperbolization [ON20] to construct many pinched Hadamard manifolds in any dimension $d\geqslant 4$, whose isometry groups admit convex cocompact subgroups with limit set a Pontryagin sphere.
@@ Acknowledgements
We thank Thierry Barbot, Genevieve Walsh and Theodore Weisman for useful conversations. We also thank Theodore Weisman for providing the beautiful Figure 1. We thank Jiming Ma for pointing out the references [Granier] and [MaXie]. S.D. thanks François Guéritaud for the content of Remark 5.4. S.D. was supported by the Huawei Young Talents Program. G.L. was supported by the European Research Council under ERC-Consolidator Grant 614733 and by the National Research Foundation of Korea (NRF) grant funded by the Korean government (MSIT) (No 2020R1C1C1A01013667). L.M. acknowledges support by the Centre Henri Lebesgue (ANR-11-LABX-0020 LEBESGUE), ANR Géométries de Hilbert sur tout corps valué (ANR-23-CE40-0012) and ANR Groupes Opérant sur des FRactales (ANR-22-CE40-0004). L.R. acknowledges support by INDAM-GNSAGA.
@@ 2.1. Simplicial complexes
  • If $L\subseteq K$ is a subcomplex, we say that $L$ is full (or induced) if whenever $d+1$ vertices span a $d$-simplex in $K$, they also span a $d$-simplex in $L$. In particular, note that a full subcomplex of a flag complex is flag, and that if $H$ is a full subcomplex of $L$ and $L$ is a full subcomplex of $K$, then $H$ is a full subcomplex of $K$.
@@
    $\operatorname{pcd}(K)=\max_{\sigma}\,\,\max\left\{n\,\left|\,\,\overline{H}^{n}(K\smallsetminus\sigma)\neq 0\right.\right\},$
    where $\overline{H}^{n}$ denotes reduced cohomology and $\sigma$ is a possibly empty subcomplex of $K$; see [DS21].
@@ 2.2. The Pontryagin sphere
The Pontryagin sphere is a $2$-dimensional continuum (i.e., connected compact metrizable space), which arises as the limit of a certain inverse system of closed surfaces of positive genus; see [JA91, FI03, ZA10, SW20, SW20sp] and Figure 1. The name “sphere” can be misleading here. Indeed, not only is this space not a manifold, but also none of its open subsets can be embedded in $\mathbb{R}^{2}$, except the empty one. Note that Pontryagin also defined an infinite family of $2$-dimensional continua called Pontryagin surfaces, which are also obtained as inverse limits of $2$-dimensional complexes but are not manifolds. However, the Pontryagin sphere is not one of the Pontryagin surfaces: while the former embeds in $\mathbb{R}^{3}$ and has rational cohomological dimension $2$, the latter do not, see [DR05, Example 1.9].
@@
Here are some more details about the construction of the Pontryagin sphere. Let $\{f_{n}:S_{n}\to S_{n-1}\}_{n\in\mathbb{N}}$ be an inverse system of closed connected orientable surfaces and continuous maps between them. Its inverse limit is called a tree of surfaces if the bonding maps $f_{n}$ satisfy some technical conditions, which roughly speaking say that $S_{n}$ is obtained from $S_{n-1}$ by performing a connected sum with some closed orientable surfaces, in such a way that the portion of $S_{n-1}$ which is engaged in the connected sum becomes dense eventually. The reader should imagine starting with a surface $S_{0}$ and performing an infinite sequence of connected sums with surfaces along smaller and smaller disks that become more and more densely packed. If all surfaces involved are $2$-spheres, then the limit is a $2$-sphere. If finitely many surfaces have positive genus, then the limit is a closed surfaces of positive genus. If infinitely many surfaces have positive genus, then the limit is the Pontryagin sphere. This space feels like a $2$-sphere with infinitely many handles attached, but since we are attaching handles all over the place, the result is a fractal space rather than a surface of infinite type, see Figure 1.
In our context, an inverse system as above arises as follows. Let $X$ be a CAT(0) cube complex in which the link of every vertex is a flag triangulation of a closed connected orientable surface of positive genus. Choose a base point $p$ in the interior of a cube and consider a family of closed metric balls $B(p,r_{n})$ of radius $r_{n}>0$ centered at $p$. Assume for simplicity that $p$ and $r_{n}$ are chosen generically, so that no ball contains a vertex on its boundary. If $B(p,r_{n})$ does not contain any vertex, then $\partial B(p,r_{n})$ is just the space of directions at $p$ and is homeomorphic to a $2$-sphere. More generally, as observed in [DJ91, §3.d], $\partial B(p,r_{n})$ is homeomorphic to the surface obtained from $\partial B(p,r_{n-1})$ by taking a connected sum with the surfaces $\operatorname{lk}(v,X)$, where $v$ ranges over the vertices contained in $B(p,r_{n})\smallsetminus B(p,r_{n-1})$. We obtain an inverse system of surfaces, in which the bonding maps are induced by the nearest point projections $B(p,r_{n})\to B(p,r_{n-1})$. Note that the collection of geodesic rays issuing from the base point $p$ and hitting a vertex is dense in the space of directions at $p$, so the connected sums involved in the inverse limit engage a dense portion of the space of directions at $p$. The visual boundary $\partial_{\infty}X$ of $X$ is the limit of this inverse system, and is homeomorphic to the Pontryagin sphere.
@@ 2.3. Right-angled Coxeter groups
$W_{K}=\langle s\in K^{(0)}\mid s^{2}=1,[s,t]=1\textrm{ iff }(s,t)\in K^{(1)}\rangle.$
The RACGs considered in this paper arise concretely as groups generated by reflections in the codimension-$1$ faces of certain polytopes in real hyperbolic $d$-space $\mathbb{H}^{d}$. More generally, a RACG $W_{K}$ can be regarded as a group generated by reflections in the hyperplanes of the associated Davis complex. This is a CAT(0) cube complex whose $1$-skeleton may be identified with the Cayley graph of $W_{K}$ and where the link of each vertex is isomorphic to the flag complex $K$. For background on Coxeter groups the reader is referred to [DA08].
@@
We now collect some well-known properties of $W_{K}$ that can be read off of its nerve for future reference. Given a full subcomplex $L\subseteq K$, the subgroup $\langle L^{(0)}\rangle$ is called the special subgroup defined by $L$.
@@
The properties of special subgroups in (1.) are classical, see e.g. [DA08, §4.1]. Item (2.) is the right-angled case of Moussong’s theorem [moussong], see also [DA08, Corollary 12.6.3]. The equality in (3.) follows from [BM91, Corollary 1.4.(e)] and [DA08, Corollary 8.5.5]. ∎
In many cases of interest, the topology of the nerve determines more than just the dimension of the Gromov boundary. For instance, we will use the following result from [FI03].
@@ 2.4. Convex cocompactness
Let $\Gamma$ be a discrete subgroup of $\mathrm{Isom}(\mathbb{H}^{d})$. The limit set of $\Gamma$ is the closed subset of $\mathbb{S}^{d-1}=\partial_{\infty}\mathbb{H}^{d}$ consisting of all accumulation points of an orbit of $\Gamma$ (this does not depend on the orbit). We say that $\Gamma$ is convex cocompact if $\Gamma$ acts cocompactly on the convex hull of its limit set.
@@
Part (1.) is [SW01, Theorem 12]. Part (2.) follows from the fact that an $H$-orbit is quasiconvex in the corresponding $\Gamma$-orbit, together with the characterization of convex cocompactness in terms of quasiconvexity of orbits from [SW01, Main Theorem]. ∎

### v3

@@ 1. Introduction
For any Gromov hyperbolic group $G$, the Gromov boundary $\partial_{\infty}G$ of $G$ is a compact metrizable space [28]. A natural problem is to characterize which spaces can arise as boundaries of Gromov hyperbolic groups (see [31, Question A]). When $\partial_{\infty}G$ is $1$-dimensional, if $G$ does not split over virtually cyclic subgroups, then $\partial_{\infty}G$ can be either $S^{1}$, the Sierpiński carpet, or the Menger curve (see [31, Theorem 4]). Moreover, the Menger curve is the generic case, see [11]. In dimensions at least $2$, no such characterization is available. Higher-dimensional spheres and Sierpiński compacta are realized by the fundamental group of real hyperbolic manifolds, respectively closed or with totally geodesic boundary. Other spaces known to be realized include certain Menger compacta, the Pontryagin surfaces, and many trees of manifolds, including the Pontryagin sphere (see [22, 26, 24, 43, 47]).
Another interesting question concerns which Gromov hyperbolic groups can be realized as discrete subgroups $\Gamma<\mathrm{Isom}(\mathbb{H}^{n})$ for some $n$, or which spaces can occur as the limit set of such groups $\Gamma$. The limit set $\Lambda_{\Gamma}$ of $\Gamma$ is a fundamental object in the study of Kleinian groups, and there is a rich theory relating the Hausdorff dimension of $\Lambda_{\Gamma}$ to the critical exponent of $\Gamma$, see [33]. In particular, when $\Gamma$ is convex cocompact, these invariants coincide, see [44]. Moreover, in this case, the abstract group $\Gamma$ is Gromov hyperbolic, and $\Lambda_{\Gamma}$ is $\Gamma$-equivariantly homeomorphic to $\partial_{\infty}\Gamma$. However, very few topological spaces are currently known to occur as limit sets. For instance, it remains an open question whether there exists a convex cocompact subgroup $\Gamma<\mathrm{Isom}(\mathbb{H}^{n})$ whose limit set is a Čech cohomology sphere other than the standard sphere (see [32, Question 9.5] and [7].)
We now turn to a more detailed discussion of the main objectives of this paper. The Sierpiński carpet is the fractal curve obtained from a square by removing smaller and smaller squares, similarly to the way one obtains the Cantor set from an interval. In particular, the Sierpiński carpet has a well-defined collection of boundary squares. The Pontryagin sphere is obtained from the Sierpiński carpet by identifying the opposite sides of each boundary square as one would glue the edges of a square to form a torus. This results in a $2$-dimensional compact metrizable space without non-empty open subsets embeddable in $\mathbb{R}^{2}$; see Figure 1. The Pontryagin sphere can also be described as the limit of a certain inverse system of closed connected orientable surfaces of positive genus; see [29, 26, 51, 47, 48] and §2.2.
The Pontryagin sphere is known to appear as the Gromov boundary of many Gromov hyperbolic groups, see for instance [43, Remark 4.4] and [30, 51], and also more recent work [25, 10] where infinitely many quasi-isometry classes of such groups are constructed. Here we consider the problem of realizing the Pontryagin sphere as the limit set of concrete reflection groups in the isometry group $\mathrm{Isom}(\mathbb{H}^{d})$ of a real hyperbolic space $\mathbb{H}^{d}$.
@@
The starting point of our work is the following. If $W$ is a Gromov hyperbolic right-angled Coxeter group (RACG) whose nerve is a closed connected orientable surface of positive genus then the Gromov boundary $\partial_{\infty}W$ of $W$ is a Pontryagin sphere; see [26]. Hence, the upshot of this paper is to find convex cocompact actions of two such Coxeter groups on $\mathbb{H}^{4}$ and $\mathbb{H}^{6}$, respectively.
@@
The Menger curve is a fractal curve that is obtained from a cube by removing smaller and smaller cubes, similarly to the way one obtains the Sierpiński carpet from a square and the Cantor set from an interval. It can be characterized as the unique compact, metrizable, connected, locally connected, $1$-dimensional space without local cut points and without non-empty open subsets embeddable in $\mathbb{R}^{2}$; see [1, 2]. According to [35], the Pontryagin sphere has a dense collection of Menger curves, so it is natural to look for Menger curves that are limit sets of quasiconvex subgroups of $\Gamma^{4}$ or $\Gamma^{6}$. In Theorem 5.3 we show that both $\Gamma^{4}$ and $\Gamma^{6}$ have convex cocompact subgroups with limit set homeomorphic to the Menger curve. These subgroups are obtained by forgetting one (or more, suitably chosen) reflections. Previous examples of convex cocompact subgroups of $\mathrm{Isom}(\mathbb{H}^{4})$ with limit set a Menger curve were constructed by Bourdon [8], see Remark 5.4.
@@
Note that the analogous statement for Gromov hyperbolic groups with boundary a $2$-sphere, namely, that every group of the latter form possesses a quasiconvex subgroup with boundary a Sierpiński carpet, would follow from the Cannon conjecture, but even this implication requires non-trivial results about hyperbolic $3$-manifolds, including several of the ingredients that went into the solution to the virtual Haken conjecture [6].
@@
We conclude with a few comments on realizing the Pontryagin sphere as a limit set in other spaces. By embedding $\mathbb{H}^{4}$ as a totally geodesic subspace in higher-dimensional hyperbolic spaces, one can realize $\Gamma^{4}$ as a convex cocompact subgroup of $\mathrm{Isom}(\mathbb{H}^{d})$ for each $d\geqslant 4$. Similarly, since $\mathbb{H}^{4}$ embeds as a totally geodesic subspace in the complex hyperbolic spaces $\mathbb{H}^{d}_{\mathbb{C}}$ for $d\geqslant 4$, in all quaternionic hyperbolic spaces, and in the octonionic hyperbolic plane, all these symmetric spaces have convex cocompact groups of isometries with limit set a Pontryagin sphere. This leaves us with $\mathbb{H}_{\mathbb{C}}^{2}$ and $\mathbb{H}_{\mathbb{C}}^{3}$. Jeffrey Danciger and Theodore Weisman have communicated to us that they can show that the deformations of certain cusped hyperbolic $3$-orbifold groups inside $\mathrm{Isom}(\mathbb{H}_{\mathbb{C}}^{3})$ constructed in [42] descend to convex cocompact representations of suitable quotients with limit set a Pontryagin sphere. On the other hand, it is known for instance that the limit set of a convex cocompact subgroup of $\mathrm{Isom}(\mathbb{H}_{\mathbb{C}}^{2})$ cannot have topological dimension $2$, see [34, Theorem 17]. In particular, the limit set of such a group cannot be a Pontryagin sphere. To summarize, there is a complete list of the negatively curved symmetric spaces whose isometry group does not contain a convex cocompact subgroup with limit set a Pontryagin sphere; these are precisely $\mathbb{H}^{2}$, $\mathbb{H}^{3}$, and $\mathbb{H}_{\mathbb{C}}^{2}$. We also note that Granier [27] constructed convex cocompact subgroups of $\mathrm{Isom}(\mathbb{H}_{\mathbb{C}}^{2})$ with limit set a Menger curve (see also recent related work of Ma–Xie [36]). Outside of the world of symmetric spaces, one can use Ontaneda’s Riemannian hyperbolization [41] to construct many pinched Hadamard manifolds in any dimension $d\geqslant 4$, whose isometry groups admit convex cocompact subgroups with limit set a Pontryagin sphere. We also remark that there exist Gromov hyperbolic groups with Gromov boundary a Pontryagin sphere, which do not admit any proper action on a real or complex hyperbolic space [37].
@@ Acknowledgements
We thank Thierry Barbot, Nikolay Bogachev, Bena Tshishiku, Genevieve Walsh, and Theodore Weisman for useful conversations. We thank Ara Bamsajian, Jiming Ma, and Subhadip Dey for pointing us to relevant references. We also thank Theodore Weisman for providing the beautiful Figure 1. S.D. thanks François Guéritaud for the content of Remark 5.4. Finally, we would like to thank the referee for carefully reading the paper and for suggesting several valuable improvements.
S.D. was supported by the Huawei Young Talents Program. G.L. was supported by the European Research Council under ERC-Consolidator Grant 614733 and by the National Research Foundation of Korea (NRF) grant funded by the Korean government (MSIT) (No 2020R1C1C1A01013667). L.M. acknowledges support by the Centre Henri Lebesgue (ANR-11-LABX-0020 LEBESGUE), ANR Géométries de Hilbert sur tout corps valué (ANR-23-CE40-0012) and ANR Groupes Opérant sur des FRactales (ANR-22-CE40-0004). L.R. acknowledges support by INDAM-GNSAGA.
@@ 2.1. Simplicial complexes
  • If $L\subseteq K$ is a subcomplex, we say that $L$ is full (or induced) if whenever $d+1$ vertices of $L$ span a $d$-simplex in $K$, they also span a $d$-simplex in $L$. In particular, note that a full subcomplex of a flag complex is flag, and that if $H$ is a full subcomplex of $L$ and $L$ is a full subcomplex of $K$, then $H$ is a full subcomplex of $K$.
@@
[…]
    where $\overline{H}^{n}$ denotes reduced cohomology and $\sigma$ is a (possibly empty) simplex of $K$; see [16].
@@ 2.2. The Pontryagin sphere
The Pontryagin sphere is a $2$-dimensional continuum (i.e., connected compact metrizable space), which arises as the limit of a certain inverse system of closed surfaces of positive genus, see Figure 1. For more details and background about the material presented in this section, we refer the reader to [29, 26, 51, 47, 48]. The name “sphere” can be misleading here. Indeed, not only is this space not a manifold, but also none of its nonempty open subsets can be embedded in $\mathbb{R}^{2}$. Note that Pontryagin also defined an infinite family of $2$-dimensional continua called Pontryagin surfaces, which are also obtained as inverse limits of $2$-dimensional complexes but are not manifolds. However, the Pontryagin sphere is not one of the Pontryagin surfaces: while the former embeds in $\mathbb{R}^{3}$ and has rational cohomological dimension $2$, the latter do not, see [23, Example 1.9].
@@
Here are some more details about the construction of the Pontryagin sphere. Let $\{f_{n}:S_{n}\to S_{n-1}\}_{n\in\mathbb{N}}$ be an inverse system of closed connected orientable surfaces and continuous maps between them. Its inverse limit is called a tree of surfaces if the bonding maps $f_{n}$ satisfy some technical conditions, which roughly speaking say that $S_{n}$ is obtained from $S_{n-1}$ by performing a connected sum with some closed orientable surfaces, in such a way that the portion of $S_{n-1}$ which is engaged in the connected sum becomes dense eventually. The reader should imagine starting with a surface $S_{0}$ and performing an infinite sequence of connected sums with surfaces along smaller and smaller disks that become more and more densely packed. If all surfaces involved are $2$-spheres, then the limit is a $2$-sphere. If finitely many surfaces have positive genus, then the limit is a closed surface of positive genus. If infinitely many surfaces have positive genus, then the limit is the Pontryagin sphere. Heuristically, this space feels like a $2$-sphere with infinitely many handles attached, but since we are attaching handles all over the place, the result is a fractal space rather than a surface of infinite type, see Figure 1.
In our context, an inverse system as above arises as follows. Let $X$ be a CAT(0) cube complex in which the link of every vertex is a flag triangulation of a closed connected orientable surface of positive genus. Choose a basepoint $p$ in the interior of a cube and consider a family of closed metric balls $B(p,r_{n})$ of radius $r_{n}>0$ centered at $p$. Assume for simplicity that $p$ and $r_{n}$ are chosen generically, so that no ball contains a vertex on its boundary. If $B(p,r_{n})$ does not contain any vertex, then $\partial B(p,r_{n})$ is just the space of directions at $p$ and is homeomorphic to a $2$-sphere. More generally, as observed in [17, §3.d], $\partial B(p,r_{n})$ is homeomorphic to the surface obtained from $\partial B(p,r_{n-1})$ by taking a connected sum with the surfaces $\operatorname{lk}(v,X)$, where $v$ ranges over the vertices contained in $B(p,r_{n})\smallsetminus B(p,r_{n-1})$. We obtain an inverse system of surfaces, in which the bonding maps are induced by the nearest-point projections $B(p,r_{n})\to B(p,r_{n-1})$. Note that the collection of geodesic rays issuing from the base point $p$ and hitting a vertex is dense in the space of directions at $p$, so the connected sums involved in the inverse limit engage a dense portion of the space of directions at $p$. The visual boundary $\partial_{\infty}X$ of $X$ is the limit of this inverse system, and is homeomorphic to the Pontryagin sphere.
@@ 2.3. Right-angled Coxeter groups
$W_{K}=\langle s\in K^{(0)}\mid s^{2}=1,\;[s,t]=1\textrm{ for all edges }(s,t)\textrm{ of }K\rangle.$
The RACGs considered in this paper arise concretely as groups generated by reflections in the codimension-$1$ faces of certain polytopes in real hyperbolic $d$-space $\mathbb{H}^{d}$. More generally, a RACG $W_{K}$ can be regarded as a group generated by reflections in the hyperplanes of the associated Davis complex. This is a CAT(0) cube complex whose $1$-skeleton may be identified with the Cayley graph of $W_{K}$ and where the link of each vertex is isomorphic to the flag complex $K$. For background on Coxeter groups the reader is referred to [19].
@@
We now collect some well-known properties of $W_{K}$ that can be read off of its nerve for future reference. Given a full subcomplex $L\subseteq K$, the subgroup $\langle L^{(0)}\rangle$ generated by the vertices of $L$ is called the special subgroup defined by $L$.
@@
The properties of special subgroups in (1.) are classical, see e.g. [19, §4.1]. Item (2.) is the right-angled case of Moussong’s theorem [40], see also [19, Corollary 12.6.3]. The equality in (3.) follows from [7, Corollary 1.4.(e)] and [19, Corollary 8.5.5]. ∎
In many cases of interest, the topology of the nerve determines more than just the dimension of the Gromov boundary. For instance, we will use the following result from [26].
@@ 2.4. Convex cocompactness
Let $\Gamma$ be a discrete subgroup of $\mathrm{Isom}(\mathbb{H}^{d})$. The limit set of $\Gamma$ is the closed subset of $\mathbb{S}^{d-1}=\partial_{\infty}\mathbb{H}^{d}$ consisting of all accumulation points of the orbit under $\Gamma$ of any point of $\mathbb{H}^{d}$ (this does not depend on the point). We say that $\Gamma$ is convex cocompact if $\Gamma$ acts cocompactly on the convex hull of its limit set.
@@
Part (1.) is [45, Theorem 12]. Part (2.) follows from the fact that an $H$-orbit is quasiconvex in the corresponding $\Gamma$-orbit, together with the characterization of convex cocompactness in terms of quasiconvexity of orbits from [45, Main Theorem]. ∎